### v1

# MOSAIC: Generating Consistent, Privacy-Preserving Scenes from Multiple Depth Views in Multi-Room Environments

Abstract
We introduce a novel diffusion-based approach for generating privacy-preserving digital twins of multi-room indoor environments from depth images only. Central to our approach is a novel Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC) model that explicitly considers cross-view dependencies within the same scene in the probabilistic sense. MOSAIC operates through a novel inference-time optimization that avoids error accumulation common in sequential or single-room constraint in panorama-based approaches. MOSAIC scales to complex scenes with zero extra training and provably reduces the variance during denoising processes when more overlapping views are added, leading to improved generation quality. Experiments show that MOSAIC outperforms state-of-the-art baselines on image fidelity metrics in reconstructing complex multi-room environments. Project page is available at: https://mosaic-cmubig.github.io.
Figure 1: For privacy-preserving scenarios where RGB collection is restricted, MOSAIC generates consistent RGB images from depth data captured along robot paths, guided by text prompts. These outputs further enable 3D reconstruction of multi-room environments.
1 Introduction
Autonomous scene reconstruction is a crucial capability in robotics and computer vision [21, 22, 14, 1], with applications spanning virtual reality, architectural design, and AI-driven simulation. However, existing multi-view 3D reconstruction methods rely heavily on capturing RGB images, which poses significant privacy risks in sensitive environments such as hospitals, factories, elder care facilities, and private residences. Capturing RGB data in these settings can inadvertently expose personal information [28, 43], limiting the practicality of generative AI solutions for real-world deployments where visual privacy must be preserved.
A promising solution for privacy-sensitive environments involves deploying mobile robots that collect only geometric data, such as depth images: this strategy preserves the scene’s structural layout while preventing the capture of sensitive texture information. In order to create digital replicas of these scenes, we need algorithms that are capable of generating high quality, multiview-consistent sequential images that align with ground truth geometry, while maintaining visual coherence across viewpoints.
Various approaches for scene level multiview image generation have been proposed, with autoregressive generation techniques [12, 7, 41, 40] representing the current mainstream trend. These approaches leverage powerful inpainting models to iteratively render unseen parts of the scene from sequential camera perspectives. However, these approaches exhibit style drift, where images generated for initial viewpoints often differ stylistically from those produced when revisiting the same location from different angles, creating visual inconsistencies. Moreover, warp-and-inpainting methods frequently encounter depth misalignment issues, leading to compounding errors over sequential generations [32, 20]. Unlike prior methods that perform diffusion synchronization tasks [2, 17, 8, 19, 15] in constrained settings (limited viewpoint variations), our task requires coherence between viewpoints based on arbitrary trajectory features, extensive perspective changes, and spatial discontinuities. This fundamental distinction makes existing methods ineffective for complex multi-room environments with widely-varying camera positions.
To address these limitations, we introduce Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC), a novel approach that achieves multi-view consistent RGB image generation from depth data, without compromising privacy in scene level, and while remaining robust to arbitrary egocentric perspective changes across multi-room environments. Rather than requiring extensive training on arbitrary trajectories, MOSAIC operates through a novel multiview inference-time optimization that ensures cross-view consistency during the diffusion process. By minimizing projection differences between overlapping views at each denoising step, our method guarantees visual coherence across the entire scene. A key advantage of MOSAIC is its ability to reduce variance and improve image-depth alignment as more overlapping views are added—the opposite behavior to current autoregressive approaches, which accumulate errors over time. Moreover, we propose a depth-weighted optimization, which naturally prioritizes information from optimal viewpoints, and a pixel-space refinement mechanism which ensures that latent-space consistency translates to visual harmony in the final output. Importantly, MOSAIC scales effortlessly to complex multi-room environments without additional training, making it practical for complex real-world applications. Our experiments demonstrate superior performance over existing methods across multiple metrics, establishing a new state-of-the-art for scene level multi-view image generation with geometric priors.
Our contributions are summarized as follows:
  1. 1.
    We propose MOSAIC, a novel diffusion-based model for multi-view-consistent image generation via explicit modeling of cross-view dependencies.
  2. 2.
    We propose a novel multi-view inference-time optimization as a practical sampling strategy that scales to complex scenes and more depth views with no extra training.
  3. 3.
    We show theoretically that with more overlapping depth views, MOSAIC reduces the variance of generation process, mitigating error accumulations in current autoregressive multi-view image generation pipelines.
  4. 4.
    We introduce depth-weighted projection loss and pixel-space refinement to enforce visual coherence. Extensive experiments demonstrate that MOSAIC generalizes across diverse viewpoints, significantly outperforming state-of-the-art methods in generation quality.
2 Related Work
Diffusion-Guided Scene-Level Multi-view Generation. Recent advances in 2D diffusion-based generative models [31, 10, 29, 42] have catalyzed significant progress in scene-level multiview image generation—particularly in the use of autoregressive generation techniques that leverage warping and inpainting mechanisms for, e.g., scene completion [32, 18, 30]. These mechanisms can introduce style drift errors where images in later iterations differ stylistically from the initial ones, leading to artifacts and distortions. Alternative approaches attempt to maintain style consistency by generating panoramic representations using off-the-shelf diffusion models [32, 18, 30, 36]. While these methods achieve internal coherence, they cannot handle multiple rooms with complex layouts or flexible camera trajectories, limiting their practicality in the real world. Recent video diffusion models demonstrate promising capabilities for generating temporally consistent image sequences  [4, 11, 39, 44]. However, they do not apply to our task due to the inability to incorporate depth conditions and limited tolerance against perspective changes. Our method, on the other hand, explicitly models cross-view consistency and generalizes to arbitrary viewpoints without imposing layout constraints.
Diffusion Synchronization. Several works have explored synchronization techniques for diffusion models [2, 17, 8, 19, 15] to achieve consistency across multiple outputs. However, existing approaches primarily address specialized cases: panoramic view generation [2, 17, 15], orthogonal transformations [8], or object-level consistency with evenly distributed camera positions and well established UV map [15, 19]. Our work addresses a fundamentally different challenge: maintaining consistent image generation across scene-level camera trajectories with arbitrary viewpoints and significant perspective variations. In this complex setting, naive averaging of denoising features fails to preserve cross-view consistency due to the non-orthogonal nature of the transformations and the extensive variance in viewpoint geometry.
Diffusion Inference-Time Optimization. To circumvent the need for large-scale model fine-tuning, several approaches have explored test-time optimization with directional guidance. Diffusion-TTA [25] adapts discriminative models using pre-trained generative diffusion models, updating parameters through gradient backpropagation at inference time. Other methods [23, 34, 24] compute directional controls for single-instance generation (music, images) by calculating directional losses during diffusion and back-propagating to update denoising features. While these methods focus on intra-level control with easily quantifiable directions, our approach is the first to propose multi-channel test-time optimization for cross-view control using self-provided directional guidance. This novel approach enables us to maintain consistency across arbitrary viewpoints without requiring additional training, addressing the fundamental limitations of existing methods when applied to complex multi-room environments.
Figure 2: MOSAIC Overview. MOSAIC implements a multi-channel denoising framework, as illustrated in (a). We formulate the training objective and address it through projection loss with inference-time optimization (b). In the final denoising phase, a pixel space refinement loss enforces precise RGB consistency constraints (c).
3 Problem Definition
We seek to acquire digital replicas of real-world, multi-room indoor environments in a privacy-preserving manner. To safeguard privacy, we refrain from collecting sensitive real RGB data and capture only geometric structures instead. From these structures, we generate synthetic RGB data—producing complex, photorealistic scenes that closely align with reality while ensuring no actual environmental details are directly revealed. In this work, we assume that the geometric structures are collected as multiple depth images $\{d^{[1]},\dots,d^{[N]}\}$ by mobile robots deployed to real-world scenes. To fully cover multi-room environments, the mobile robot plans proper camera pose sequences to ensure overlapping views. Namely, for each $d^{[i]}$, there must exist some area that is also covered by at least one other view $d^{[j]}$ $(j\neq i)$. From $\{d^{[i]}\}_{i\in[N]}$, we are interested in generating corresponding RGB images $\{x^{[i]}\}_{i\in[N]}$ that are multi-view consistent: overlapping depth views must lead to consistent RGB outputs where they overlap in ${\bar{x}}$. Namely, for each pixel $p$ in ${\bar{x}}$ that is covered by multiple depth views, i.e., $I_{p}:=\{i\in[N]\mid p\in d^{[i]}\}$, the RGB output should agree:
$x^{[i]}[p]=x^{[j]}[p]={\bar{x}}[p],\forall i,j\in I_{p},i\neq j$ (1)
Then the generated RGB would form a complete 3D multi-room scene ${\bar{x}}$ once warped by the cameras poses $\{T^{[i]}\}_{i\in[N]}$ (e.g., ${\bar{x}}=\sum_{i\in[N]}T^{[i]}x^{[i]}$).
4 Method: MOSAIC
4.1 Preliminaries
Denoising Diffusion Implicit Models (DDIM) [31] train a generative model $p_{\theta}(z_{0})$ to approximate a data distribution $q(z_{0})$ given samples $z\in\mathcal{Z}$ from $q$. DDIM considers the following non-Markovian inference model:
$q_{\sigma}\left(z_{1:T}\mid z_{0}\right):=q_{\sigma}\left(z_{T}\mid z_{0}\right)\prod_{t=2}^{T}q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ (2)
where $\sigma\in\mathbb{R}_{\geq 0}^{T}$ is a real vector and $z_{1:T}$ are latent variables in $\mathcal{Z}$. The forward process $q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ is parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{\top}$. To approximate $q(z_{0})$, DDIM learns a generative process
$\quad p_{\theta}\left(z_{0:T}\right):=p_{\theta}\left(z_{T}\right)\prod_{t=1}^{T}p_{\theta}^{(t)}\left(z_{t-1}\mid z_{t}\right).$ (3)
Starting from a prior $p_{\theta}(z_{T})=\mathcal{N}(0,I)$, $z_{t}$ is sampled by $p_{\theta}^{(t)}(z_{t-1}\mid z_{t})=q_{\sigma}\left(z_{t-1}\mid z_{t},f_{\theta}^{(t)}(z_{t})\right)$, where
$f_{\theta}^{(t)}\left(z_{t}\right):=\left(z_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}^{(t)}\left(z_{t}\right)\right)/\sqrt{\alpha_{t}}$ (4)
predicts $z_{0}$ with a noise prediction model $\epsilon_{\theta}^{(t)}$. Learning is performed by optimizing the standard variational objective
$J_{\sigma}\left(\epsilon_{\theta}\right):=\mathbb{E}_{z_{0:T}\sim q_{\sigma}}\left[\log q_{\sigma}\left(z_{1:T}\mid z_{0}\right)-\log p_{\theta}\left(z_{0:T}\right)\right]$ (5)
In this work, we generate RGB images from depth views $d$, via ControlNet [42]: $\epsilon_{\theta}^{(t)}(z_{t},d)$. In addition, we sample latent RGB representations $z$ from $p_{\theta}$ which can later be decoded into full images via a pre-trained VAE decoder [16].
In short, the generation process we consider is
$z_{0}^{[i]}\sim p_{\theta}(z_{0},d^{[i]}),~{}x^{[i]}=g(z_{0}^{[i]})~{}\text{for}~{}i\in[N].$ (6)
For clarity, we drop the depth dependency $d^{[i]}$ in notations in the rest of this paper. While the generation of a single image $x_{i}$ can be readily solved via depth-conditioned diffusion models [42], the generation of complete scenes from multiple depth images remains unsolved.
4.2 MOSAIC Formulation
Intuitively, the sampling procedure in Eq. 6 would generate inconsistent RGB views because the generative process $p_{\theta}$ is trained with independent samples $z_{0}\sim q(z_{0})$. In our case, however, the samples $z_{0}^{[1:N]}$ are indeed dependent since they are taken from the same complete scene ${\bar{z}_{0}}$. To fundamentally address the consistency issue, our key insight is to explicitly model such dependency by incorporating extra projection-based conditionals $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$ in the forward procedures. In practice, given ${\bar{z}_{0}}$, $z_{0}^{[i]}$ can be collected by simply sampling a camera pose with projection function $\pi^{[i]}$ and setting $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}})$; subsequent latents $z_{1:T}^{[i]}$ can be sampled as usual. The updated forward procedure for $N$ depth views from the same scene ${\bar{z}_{0}}$ can be rewritten from Eq. Eq. 2 as:
$q_{\sigma}(z_{0:T}^{[1:N]}\mid{\bar{z}_{0}})=\prod_{i\in[N]}q(z_{0}^{[i]}\mid{\bar{z}_{0}})q_{\sigma}(z_{1:T}^{[i]}\mid z_{0}^{[i]}).$ (7)
The corresponding generative process can be extended to
$p_{\theta,\phi}(z_{0:T}^{[1:N]},{\bar{z}_{0}})=p_{\phi}({\bar{z}_{0}}\mid z_{0}^{[1:N]})\prod_{i\in[N]}p_{\theta}(z_{0:T}^{[i]}),$ (8)
where $\phi$ parameterizes the reverse process of image projection $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$. We name the model $p_{\theta,\phi}$ as Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC), which can be learned by minimizing:
$J_{\sigma,\mathrm{MOSAIC}}\left(\epsilon_{\theta},\phi\right):=$
$\mathbb{E}_{(z_{0:T}^{[1:N]},{\bar{z}_{0}})\sim q_{\sigma}}\left[\log q_{\sigma}\left(z_{0:T}^{[1:N]}\mid{\bar{z}_{0}}\right)-\log p_{\theta,\phi}\left(z_{0:T}^{[1,N]},{\bar{z}_{0}}\right)\right].$ (9)
4.3 Sampling from MOSAIC with Multi-Channel Inference-Time Optimization
While MOSAIC captures the dependency between depth views from the same scene, training directly using Eq. 9 is infeasible for two reasons: (a) optimizing the expectation in Eq. 9 via stochastic batches can be computationally intractable, since sufficient combinations of $N$ views are required even from a single scene. Meanwhile, $N$ should be sufficiently large to effectively cover multi-room environments; and (b) $p_{\theta,\phi}$ trained with some fixed $N$ cannot easily generalize to various depth view numbers which further limits the applicability. Hence, we desire a tractable strategy to sample from MOSAIC that is also compatible with an arbitrary number of depth views $N$. Factorizing $q_{\sigma}$ using Eq. 7 and $p_{\theta,\phi}$ using Eq. 8, we obtain:
$J_{\sigma,\mathrm{MOSAIC}}\left(\epsilon_{\theta},\phi\right)$
[…]
$~{}~{}~{}~{}~{}-\mathbb{E}_{q_{\sigma}}\big{[}\log p_{\phi}({\bar{z}_{0}}\mid z_{0}^{[1:N]})\big{]}.$ (10)
Note that $p_{\phi}({\bar{z}_{0}}\mid z_{0}^{[1:N]})$ essentially approximates the posterior of the projection operation in forward process and evaluates how well the original scene ${\bar{z}_{0}}$ is reconstructed from separate views with some parameterization $\phi$. Hence, we can reasonably set $p_{\phi}$ to have an inverse exponential dependency on the total re-projection error:
$p_{\phi}({\bar{z}_{0}}\mid z_{0}^{[1:N]})\propto\exp(-\textstyle\sum_{i}\|{\bar{z}_{0}}-f_{\phi}^{[i]}(z_{0}^{[i]})\|)$ (11)
where $f_{\phi}^{[i]}$ projects individual views back to the original and error is considered within the back projected region. Hence the second term in Eq. 10 can be written as $\mathbb{E}_{q_{\sigma}}\big{[}\textstyle\sum_{i\in[N]}\|{\bar{z}_{0}}-f_{\phi}^{[i]}(z_{0}{[i]})\|\big{]}$. Since ${\bar{z}_{0}}$ is in general unavailable, we can instead minimize the total projected cross-view error:
$J_{\mathrm{Proj}}(\phi)=\mathbb{E}_{q_{\sigma}}\big{[}\textstyle\sum_{i,j}\|f_{\phi}^{[i]}(z_{0}^{[i]})-f_{\phi}^{[j]}(z_{0}^{[j]})\|\big{]}\geq 0$ (12)
Hence, Eq. 10 is equivalent to
$J_{\sigma,\mathrm{MOSAIC}}\left(\epsilon_{\theta},\phi\right)\equiv\textstyle\sum_{i\in[N]}J_{\sigma}^{[i]}(\epsilon_{\theta})+J_{\mathrm{Proj}}(\phi).$ (13)
Although it is impractical to directly train $p_{\theta,\phi}$ using Eq. 13, it is possible to approximate the samples from $p_{\theta,\phi}$ by fine-tuning the output of pre-trained models $p_{\theta}$, e.g., Latent Diffusion Model (LDM) [29]. We observe that for MOSAIC, $p_{\theta,\phi}$ fits trajectories $z_{1:T}^{[1]},\dots,z_{1:T}^{[N]}$ that are mutually dependent at each step $t=1,\dots,T$ through $z_{0}^{1:N}$, while such dependency is missing from LDM $p_{\theta}$. Furthermore, this dependency is exactly captured by $J_{\mathrm{Proj}}(\phi)$. If $p_{\theta,\phi}^{*}$ optimally solves $J_{\sigma,\mathrm{MOSAIC}}$ and $z_{0}^{[1:N]}\sim p_{\theta,\phi}^{*}$ , there must exist some ${\bar{z}_{0}}$ and corresponding projections $\pi^{[1:N]}$ such that $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}}),~{}\forall i\in[N]$. Then, $J_{\mathrm{Proj}}(\phi)$ must have been minimized to 00 with $f_{\phi^{*}}^{[i]}=(\pi^{[i]})^{-1},~{}\forall i$. In other words, samples from an ideal MOSAIC model should have zero projected cross-view error in expectation as long as $f_{\phi}^{[i]}$ exactly matches the inverse projection $(\pi^{[i]})^{-1}$. While it is generally hard to fit an inverse projection, it is unnecessary in our case because the ground truth projections $\pi^{[1:N]}$ are directly available when collecting depth conditions $d^{[1:N]}$ [^1] With mobile robots, $\pi^{[i]}$ can be computed from the camera/robot poses in the world coordinate when taking each depth view $d^{[i]}$..
Hence, we can approximately sample from MOSAIC by fine-tuning LDM (i.e., $p_{\theta}$) output at each denoising step $t-1$ by solving the following inference-time optimization through gradient decent:
$\underset{z_{t-1}^{[1:N]}}{\mathbf{min}}$ $~{}L_{\mathrm{Proj}}(z_{t-1}^{[1:N]})~{}\textbf{s.t.}~{}z_{t-1,\text{init}}^{[i]}\sim p_{\theta}^{(t)}(z_{t-1}^{[i]}\mid z_{t}^{[i]}),$ (14)
where the empirical projection loss is given by
$L_{\mathrm{Proj}}(z_{t}^{[1:N]})=\sum_{i,j}\|(\pi^{[i]})^{-1}(\hat{z}_{0}^{[i]})-(\pi^{[j]})^{-1}(\hat{z}_{0}^{[j]})\|.$ (15)
$\hat{z}_{0}^{[i]}$ is the estimated $z_{0}^{[i]}$ from $z_{t}^{[i]}$, i.e., $\hat{z}_{0}^{[i]}=f_{\theta}^{(t)}(z_{t}^{[i]})$.
Depth Weighted Projection Loss. When calculating $L_{\mathrm{Proj}}$, we observe that depth information provides a natural weighting mechanism for view contributions. For points visible from multiple views, the view with smaller depth value (i.e., closest to the camera) likely provides more accurate RGB information due to higher sampling density and reduced occlusion. To incorporate this insight, we modify our projection loss to weight view contributions based on their relative depth values:
$L_{\text{Proj}}^{\text{DW}}(z_{t}^{[1:N]})=\sum_{i,j}w_{i,j}\cdot\|(\pi^{[i]})^{-1}(\hat{z}_{0}^{[i]})-(\pi^{[j]})^{-1}(\hat{z}_{0}^{[j]})\|$ (16)
where $w_{i,j}$ weights the importance of consistency between views $i$ and $j$ based on their relative depths. To prioritize views with smaller depth values, we employ weighting scheme:
$w_{i,j}=\frac{\exp(-\alpha\cdot(\pi^{[i]})^{-1}d^{[i]})}{\exp(-\alpha\cdot(\pi^{[i]})^{-1}d^{[i]})+\exp(-\alpha\cdot(\pi^{[j]})^{-1}d^{[j]})},$ (17)
where $\alpha$ is a hyperparameter controlling the selectivity of the weighting. With this formulation, when aggregating across all views, the final RGB values for overlapping regions will naturally favor views with minimal depth values.
4.4 Pixel Space Refinement Loss
While our projection loss effectively ensures consistency in latent space, the non-linear nature of the VAE [16] decoder means that consistency in latent space does not necessarily translate to pixel-space consistency. This discrepancy is particularly pronounced when projection transformations are far from orthogonal. We address this issue through a pixel space refinement process during the final denoising stages. For each view $i$, we decode its predicted latent into pixel space: $\hat{x}^{[i]}=g(\hat{z}_{0}^{[i]})$. It is infeasible to calculate $L_{Proj}$ in pixel space, because back-propagation through multi-channel VAE decoder poses computation difficulties. We propose to warp these decoded images between views and compute representations for each view point by depth-weighted fusing the overlapped views:
$w_{ij}=\frac{\exp(-\alpha\cdot\pi_{ij}d^{[j]})}{\sum_{k=1}^{N}\exp(-\alpha\cdot\pi_{ik}d^{[k]})},~{}{x}^{[i]*}=\sum_{j=1}^{N}w_{ij}\cdot\hat{x}^{[j]}$ (18)
where $\alpha$ control the selectivity of weighting and $\pi_{ij}$ is the image warping from $j$ view to $i$ view. These optimal pixel-space representation are re-encoded to latent space: ${z}_{0}^{[i]*}=\beta\cdot f({x}^{[i]*})$. The Pixel Space Refinement Loss is defined as:
$L_{Pixel}=\sum_{i=1}^{N}\|\hat{z}_{0}^{[i]}-{z}_{0}^{[i]*}\|$ (19)
This directly addresses the latent-to-pixel mapping inconsistencies by forcing latent representations to align with those derived from optimally blended pixel-space images. We utilize the $L_{Pixel}$ at the final denoising stages.
4.5 Properties of MOSAIC
Training-free inference time scale-up. The sampling scheme in Eq. 14 essentially consists of two parts: (a) encouraging multi-view consistency via $L_{\mathrm{Proj}}$ which is defined for an arbitrary number of views $N$; and (b) progressing the reverse process by invoking LDM independently on each view. Hence, unlike full training using Eq. 9, sampling from MOSAIC is agnostic of $N$, allowing our generation process to easily scale and adapt to larger scenes with more views during inference time, with no extra training.
Variance reduction for pre-trained LDMs. Since LDMs are normally pre-trained in scale, we can roughly view them to have zero prediction error in expectation. However, as will be shown later, LDMs can generate results with varying quality (e.g., in terms of depth preservation) due to the stochasticity of the denoising process, which can cause errors to accumulate until the whole process fails eventually in the warp-inpainting approach. To this regard, one important advantage of MOSAIC is its ability to stabilize the denoising process given more overlapping views. This can be seen by analyzing the expected variance of a scene $\bar{z}_{0}$ given a varying number of views $z_{0}^{[1:N]}$ that overlap. Consider a partition $z_{0}^{[1:N]}=\{z_{0}^{[I_{1}]},z_{0}^{[I_{2}]}\}$ where $I_{1}\cup I_{2}=\{1,\dots,N\}$,
$\Sigma(\bar{z}_{0}\mid z_{0}^{[I_{1}]})$ $=\mathbb{E}[\Sigma(\bar{z}_{0}\mid z_{0}^{[I_{1}]},z_{0}^{[I_{2}]})]+\underbrace{\Sigma(\mathbb{E}[\bar{z}_{0}\mid z_{0}^{[I_{1}]},z_{0}^{[I_{2}]}])}_{\Sigma\text{~{}explained by~{}}z_{0}^{[I_{2}]}}$
$\succeq\mathbb{E}[\Sigma(\bar{z}_{0}\mid z_{0}^{[I_{1}]},z_{0}^{[I_{2}]})].$ (20)
Taking expectation over $z_{0}^{[I_{1}]}$, we have
$\mathbb{E}[\Sigma(\bar{z}_{0}\mid z_{0}^{[1:N]})]\preceq\mathbb{E}[\Sigma(\bar{z}_{0}\mid z_{0}^{[I_{1}]})].$ (21)
Namely, when conditioned on more views $z_{0}^{[I_{2}]}$, the variance of the final output $\bar{z}_{0}$ reduces in expectation by an amount positively related to the overlap between $\bar{z}_{0}$ and $z_{0}^{[I_{2}]}$ [^2] The amount reduced represents the variability of $\bar{z}_{0}$ due to changes in $z_{0}^{[I_{2}]}$, which is higher when $z_{0}^{[I_{2}]}$ covers more area in $\bar{z}_{0}$.. The same applies when $z_{0}^{[1:N]}$ is replaced by estimates from intermediate $z_{t}^{[1:N]}$, hence the variance reduction happens during the entire denoising process.
Table 1: Quantitative evaluation with Kernel Inception Distance (KID), CLIPIQA+, CLIPScore (CS) and CLIPConsistency(CC).
Method KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ CC $\uparrow$
Warp-Inpaint [20] 0.04646 0.6849 0.6517 29.93
MVDiffusion [33] 0.03640 0.7016 0.6025 29.41
SceneCraft [38] 0.07697 0.7011 0.4531 27.68
Text2Room [12] 0.08594 0.6618 0.4388 27.34
MOSAIC (ours) 0.03391 0.7166 0.6526 30.85
5 Experimental Setup
Navigation Agent Autonomous Data Collection. To obtain testing data, we deployed an indoor navigation robot to explore various indoor scenes, collecting depth information via onboard sensors while recording precise camera poses along its trajectory. This approach ensures aligned trajectory-based inputs with baseline methods [33, 38, 12, 20]. To evaluate generalizability, we gathered data across 16 HM3D [27] and 5 MP3D [6] scenes. For trajectory sample details, please refer to Appendix. To generate captions for these depth images, we visualized the collected depth data and utilized a VLM: ChatGPT-4o [5] to provide contextually relevant scene descriptions. This results in 2011 depth data, camera pose, caption pairs in total.
Baselines. We compare our approach against state-of-the-art methods of multiview image generation with geometric priors, including MVDiffusion [33], Warp-Inpaint [20], and SceneCraft [38]. Although SceneCraft shares different condition settings, where its input is the semantic map and the depth map of the bounding box, we create the layout of the bounding box according to our dataset for fair comparison. We evaluate semantic quality against Text2Room [12], a specialized method for indoor scene generation.
Evaluation metrics. The goal of multiview image generation with depth prior is to produce images that maintain consistency in overlapping regions, while preserving image quality, style coherence, and accurate geometric alignment. To evaluate image consistency, we follow MVDiffusion [33] by calculating Warped PSNR [13] and Warped Ratio. We assess image quality using Kernel Inception Distance (KID) [3] and CLIPIQA+ [35], and measure image-text alignment with CLIPScore [9] and CLIPConsistency [26].
Figure 3: Qualitative comparison with multi-view generation baselines. More results are available in the supplementary material.
6 Results
6.1 Evaluation Against Baselines
Qualitative Comparisons.
As shown in Fig. 3, MOSAIC produces more coherent and photorealistic results than existing approaches using identical depth inputs. MVDiffusion [33] lacks cross-view consistency and cannot generalize to stylistic prompts like “in Van Gogh style.” The warp-and-inpainting approach [20] generates acceptable initial frames but suffers from progressive style drift and accumulating errors during sequential generation. SceneCraft [38] fails to capture complex environment layouts and produces blurry artifacts due to its NeRF-based methodology. Text2Room [12], built on the warp-and-inpainting pipeline, exhibits noticeable patchwork-like seams from inpainting inconsistencies as well as the depth misalignment. In contrast, MOSAIC maintains consistency across viewpoints while preserving fidelity to both depth information and style prompts, effectively eliminating boundary artifacts through our depth-weighted projection mechanism.
Quantitative Comparisons.
Table 2: Cross-view consistency analysis.
Method Warped PSNR ↑ Warped Ratio ↑
GT 25.45 1.00
MVDiffusion 13.58 0.53
Warp-Inpaint 22.00 0.86
Ours Pixel 25.30 0.99
Tab. 1 presents a quantitative comparison of our proposed method against several state-of-the-art approaches, including Warp-and-Inpainting [20], MVDiffusion [33], SceneCraft [38], and Text2Room [12], on the multi-view consistent image generation task. Our method achieves the lowest KID score ($\downarrow$) of 0.0391 and highest CS ($\uparrow$) of 0.7166, CIQA of 0.6526, and CC of 30.85, outperforming other methods by a significant margin in all image quality metrics. In addition to the quality improvement, Tab. 2 demonstrates our exceptional cross-view consistency with a Warped PSNR of 25.30 and Warped Ratio of 0.99, approaching ground truth performance (25.45, 1.00) and substantially outperforming competing methods. Although MVDiffusion achieves a competitive KID score of 0.0640, its cross-view consistency is highly limited with a Warped PSNR of only 13.58. Our approach achieves outstanding performance in both image quality and cross-view consistency compared with state-of-the-art methods.
6.2 Ablation Study
Figure 4: Qualitative ablation results. Red: artifacts; orange: pixel space texture misaligned; Green: with pixel space loss, full model (MOSAIC) aligns the pixel space color.
Effect of MOSAIC Objective and Multi-Channel Test-Time Optimization. We formulate a naive ablation baseline by extending [15] to scene-level multi-view image generation, averaging each denoising latent prediction $z_{0}$ without our multi-channel inference time optimization for MOSAIC objectives. As shown in Tab. 3, this naive approach yields poor results with the highest KID score and lowest consistency metrics. Fig. 4 reveals significant issues—the second row exhibits severe blurry artifacts (red boxes) and compromised cross-view consistency. In contrast, our full model with pixel space refinement delivers exceptional performance, achieving the lowest KID score, highest consistency metrics, and superior cross-view alignment (green boxes). These results demonstrate that MOSAIC objective optimization is crucial for multi-room environments with significant viewpoint variations, where naive averaging fails to maintain both visual quality and geometric coherence.
Effect of Pixel Space Refinement. Tab. 3 shows how pixel space refinement enhances MOSAIC’s performance. This refinement yields notable improvements in both perceptual quality (lower KID) and cross-view consistency (approaching ground truth metrics). Fig. 4 illustrates this improvement between 3rd and 4th row: texture misalignments (yellow) become properly aligned (green) after refinement.
Effect of Trajectory Key Frame Selection. We evaluate our region-coverage sampling strategy against uniform trajectory sampling in Tab. 3. While image quality and consistency metrics remain comparable between approaches, text-image alignment noticeably decreases with uniform sampling. This occurs because uniform sampling frequently captures frames facing flat walls, challenging VLMs to generate precise captions. Furthermore, pretrained ControlNet models tend to hallucinate textures with such depth inputs, creating semantic misalignments. Our strategic key-frame selection effectively balances information-rich viewpoints with sufficient overlap, yielding more accurate text-conditioned generation without sacrificing consistency.
Table 3: Ablation: Image Quality and Consistency
Image Quality Image Consistency
KID $\downarrow$ CS $\uparrow$ CIQA $\uparrow$ CC $\uparrow$ PSNR $\uparrow$ Ratio $\uparrow$
w/o test-time optimization 0.06715 0.7285 0.5725 31.08 15.74 0.6185
w/o key frame selection 0.03920 0.6746 0.6733 29.25 24.67 0.9692
w/o $L_{Pixel}$ 0.04273 0.7249 0.6797 31.26 23.02 0.9045
MOSAIC (ours full) 0.03391 0.7166 0.6526 30.85 25.30 0.9940
Figure 5: Second column: Single-view generation creates artifacts (in red) contradicting ground truth depth. Third column: MOSAIC achieves superior depth alignment (in green).
Multi-Views for better RGB-Depth Alignment. Traditional scene-level generation approaches employ warp-and-inpainting strategies that suffer from error accumulation—initial RGB-depth misalignments compound through sequential generation. In contrast, our theoretical analysis in Eq.21 establishes that leveraging multi-view depth conditioning inherently reduces variance, yielding representations that more faithfully reflect ground truth geometry. Tab. 4 empirically validates this theoretical insight: as viewpoint count increases, normalized MSE between ground truth depth and generated image depth (estimated via Depth Anything [37]) demonstrates consistent decline. Fig. 5 visually confirms this phenomenon—single-view conditioning (middle column) produces hallucinated elements that misalign with ground truth geometry (red boxes), while MOSAIC’s two-view approach (right column) maintains precise geometric fidelity. Our analysis further reveals an inflection point in the quality-viewpoint curve, with text-image alignment steadily improving across viewpoints while perceptual quality peaks at 2-3 highly overlapping views. This finding suggests an optimal operating point that balances computational efficiency with generation quality.
Table 4: Assessing the effect of number of depth views, NMSE stands for normalized MSE.
NMSE $\downarrow$ CS $\uparrow$ KID $\downarrow$
single view 0.018630 0.6874 0.06123
two views 0.013453 0.6883 0.06022
three views 0.011961 0.6986 0.06709
four views 0.011269 0.7100 0.07833
7 Conclusion
We proposed MOSAIC, the first training-free multi-view consistent image generation pipeline that operates at scene level. MOSAIC handles arbitrary numbers of views along any trajectory, adapting to extensive camera viewpoint changes. It employs a depth-weighted projection loss and a pixel-space refinement process to maintain visual coherence across complex multi-room environments, making it well-suited for privacy-sensitive applications. We developed a novel multi-channel inference-time optimization procedure that minimizes cross-view projection errors, which we mathematically prove to be equivalent to the ideal learning objective. We showed theoretically and demonstrated experimentally that by intelligently fusing multi-view information, MOSAIC substantially improves alignment between generated RGB images and ground truth depth and significantly outperforms state-of-the-art methods—addressing the error accumulation issues of current autoregressive multi-view image generation pipelines.
Acknowledgements
This work was supported in part by NSF IIS-2112633.